Shift report 20030307 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Mar 07 2003 - 08:02:07 EST

  • Next message: Stephen J. Sanders: "Re: BNL workshop 6-8 March"
    ------------------------------------------------------------
    Date: Fri Mar  7 08:02:07 2003, Shift: 00:00-08:00
    Supervisor: ZY, with: 
    Log book#14 pages 50 - 50
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 2 o'clock till now. 
    2. Data runs:
      For 3deg FS 1/2A and 90deg MRS A1050,
        RUN 8242 52k trig 2 and 65k trig 3
        RUN 8243 86k trig 2 and 104k trig 3
      For 3deg 1/4A and 90deg MRS A1050,
        RUN 8244 194k trig 2 and 108k trig 3
        RUN 8245 114k trig 2 and 63k trig 3
      Now we are in 4deg 2/3A and 90deg MRS. I have voted to dump the beam.
    3. No detector problem.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Mar 07 2003 - 08:02:42 EST